Test whether the series ∞Σn=0 1/(n^5+x^3) converges uniformly or not

1
Expert's answer
2022-03-15T09:31:24-0400

The series converges for all x such that "n^5+x^3\ne 0" for "n\geq 1" .

Consider

"\left| S_n\left( x \right) -S_{n-1}\left( x \right) \right|=\left| \frac{1}{n^5+x^3} \right|"

For "x=-\left( n^5-1 \right) ^{1/3}"

"\left| S_n\left( x \right) -S_{n-1}\left( x \right) \right|=1"

The series doesn’t converge uniformly by the Cauchy criterion.